A call option on ABC Inc. has a strike price of $25, a bid price of $0.20 and ask price of $0.25. The current share price is $24.25. If the future share price of ABC could be either $22 or $35, and the current risk-free rate is 5%, what is the value of this call option?

0.5

2.54 (wrong answer)

0.2

3.75

2.90
Answer #1

CB CCCD CE CG CH CI CJ 0.9 2) d = 22/24.25 u= 35/24.25 Strike price = Spot Price = Future price = Risk free rate = $25 $24.25